# Privacy Bills of Materials: A Transparent Privacy Information Inventory for Collaborative Privacy Notice Generation in Mobile App Development

### v1

@@ 3 The PriBOM Approach
Figure 3: An overview of PriBOM in the usage scenario. The interaction between app users and UI components triggers data practices. These practices are disclosed to users through privacy notices such as the privacy policy. PriBOM helps development team create accurate privacy notices by documenting privacy information related to specific UI components.
3.1 Motivation of PriBOM
@@
Bill of Materials (BOM). To cater to increasingly complex and diverse software systems, Bill of Materials (BOM) has been introduced as the information inventory with a focus on the transparency of software materials. Proposed in the 2010s [20], SBOM has become a key strategy in response to the emerging risks in software vulnerabilities management, license compliance, and supply chain transparency [81, 60]. SBOMs record the material details of the components in developing software, providing transparency, traceability and verifiability to software building and managing process [72, 87, 58]. The US Presidential Executive Order (EO) 14028 on Improving the Nation’s Cybersecurity [^3] US Presidential Executive Order 14028 requires companies providing services to the US government to provide SBOMs. Approximately 78% of the organizations worldwide would use SBOMs by 2022 and 88% by 2023 [72, 38]. In addition to SBOM, the concept of BOM has been implemented into other forms, including HBOMs [35] and FBOMs [22, 34, 27] for hardware and firmware, DataBOMs [26] for datasets, and AIBOMs [32, 80] for AI models. Consequently, we adapt a similar design for privacy information and introduce the concept of PriBOM. PriBOM records software privacy information in a structured way and with a focus on UI, pulling different development roles to the same page. Specifically, PriBOM is indexed by the UI widget, including privacy information such as associated privacy permissions, data types, third-party libraries, and corresponding privacy notice disclosures. We will elaborate the design in Section 3.2.
@@
UI Widgets - The Pivot of Mobile Privacy Communication. The UI widgets serve as both visual elements and key components in functionality and data handling, therefore it is the pivot to synergistically connect different roles on privacy communication. First, UI components are fundamental in the software development process, which are not only visual elements but are often directly involved in the app’s functionality and data handling. The transparency regarding what data is collected through specific UI components is one of the information that users want to know [62], especially when malicious design may lead to unwanted data collection and cause harm to users [29, 33]. Moreover, UI represents the tangible interface where users interact with the application, making it a critical point that is associated with potential privacy issues. When conducting privacy-related tests, interaction with the UI widgets is also a crucial part of executing privacy checks [76, 59]. Leveraging UI to link privacy information does inherently focus less on non-UI components, which may lead to gaps in capturing practices that circumvent standard permission protocols [68]. Our choice is a deliberate trade-off to prioritize ease of use for all stakeholders over exhaustive documentation, fostering a more accessible approach to privacy management across different roles.
@@
Table 2: Design of PriBOM. One example of filled PriBOM of a real mobile application [8] is presented.
@@ 3.2 Design of PriBOM
  • [Permission] Android utilizes app permission mechanism [12] to protect access to both restricted data, e.g. Contacts, and restricted actions, e.g. taking pictures. More specifically, an app must require runtime permissions before it obtains access to additional data or performs actions that may affect the system, other apps or devices. Such requests may be involved in the widget’s behind-the-scene behaviors [79]. For example, an elliptical coordinate icon in a food delivery app is likely to be accompanied by a request to grant geographic location permissions. Such requests are raised by the Android API calls in the widget callbacks.
@@
  • [Data Type] This field records the specific data type required by corresponding Android Permission, which is also regarded as the data that the widget collects or processes. Following the grouping strategy in previous studies [56, 67] and Official Android API Documentation [10], we map the Android Permissions to specific data types. For example, “READ_CONTACTS” and “WRITE_CONTACTS” are categorized into data type “Contacts”. A more detailed discussion is in Section 4.1.
@@
  • [TPL Information] Third-party libraries (TPL) pervasively serve as reusable functional components for mobile apps to improve development efficiency [78]. For example, the in-app payment service of a shopping app may be supported by TPLs. PriBOM, therefore, records the baseline information of TPLs relevant to the widget, including name, version, latest version, publish date (current version), and publish date (latest version). PriBOM documents the publish date of the currently used and most up-to-date version of TPLs to enable assessing the currency of the version in use and informing about potential updates needed for security or functionality enhancements.
Privacy Notice Disclosure Section. In application development, a significant portion of effort is often allocated to ongoing updates rather than the outset. This principle also applies to privacy notices, given the evolving nature of legal and regulatory requirements. Consequently, PriBOM effectively manages two common forms of privacy notices, e.g., privacy policies and privacy labels, by linking widget components directly to their respective disclosures. This systematic linkage ensures that the app’s data practices are consistently aligned with legal requirements throughout the development lifecycle.
@@
  • [Privacy Label Declaration] Before publishing apps on markets such as Google Play, developers must declare their app’s privacy practices, e.g. what types of user data will be collected and how it will be handled, in a privacy nutrition label form to inform potential users. Due to the fact developers struggling to create authoritative and accurate privacy labels [52, 24, 25], including privacy labels’ data practice declarations that match the data type in PriBOM can support developers and legal teams in aligning privacy disclosure with actual application behaviors.
3.3 Benifits of PriBOM
@@
[Benefit-1] Transparency. Lacking team-level and organizational support is considered a challenge faced by developers [49, 52]. Therefore, to address the privacy ambiguity caused by the lack of team support and assistant tools, the necessity of a recording and communication tool focused on privacy information has become more prominent. PriBOM can serve as a central communication tool, bridging the informational gap between developers of different roles from various teams. By providing a unified view of privacy practices at the widget level, PriBOM can help all developers have a consistent understanding of potential privacy implications, fostering informed decision-making and cohesive development strategies. Through PriBOM, transparency can be achieved as every practitioner, regardless of their role, gains clear insights into privacy aspects.
[Benefit-2] Privacy Traceability and Trackability. PriBOM connects the frontend UI widget, the backend privacy practices in source code, and the app’s privacy notices, helping developers gain traceability and trackability on this privacy chain. Similar to terminology in software supply chain [40], both terms in the context of PriBOM with ‘where-from’ and ‘where-used’ [66] directions respectively are as follows:
  • Tracing. Given any point in this privacy chain, PriBOM enables developers to work backward to query privacy-relevant information. For example, if a user reports privacy issues in interacting with certain UI widgets, then the development team can easily track back and locate the problematic code files (Currently, a privacy breach takes an average of 287 days to completely resolve [19]).
@@ 4 Pre-fill of PriBOM
In this section, we present a pre-fill implementation of PriBOM, and describe details regarding the pipeline. Our pre-fill implementation is built upon existing tools [69, 83, 84, 2, 78, 62]. Notably, although our design tailors to Android apps, the concept of PriBOM can be easily adapted on iOS apps with corresponding modifications.
4.1 Static Analysis
Figure 4: The pipeline of our static analysis module.
@@ 4.1 Static Analysis
Following Android analysis approaches in previous icon-behavior researches [79, 88, 55], our static analysis-based pre-fill implementation consists of four steps: (1) Widget and Callback Method Extraction; (2) Call Graph Construction; (3) Android Permission Extraction; and (4) Permission-Data Mapping. Figure 4 provides an overview of the pipeline.
@@
Widget and Callback Method Extraction. In the first step, the implementation pipeline takes the Android application package (APK) file as input, and leverages JADX [18] to disassemble it. The purpose of this step is to extract widget components, associated sources such as icons, and corresponding callback methods that respond to UI interactions. Icons (such as PNG or JPG files) are commonly bound with UI widgets in the layout configuration file. For example, XML attribute android: background is used to set a background image for a widget component. Therefore, we parse the layout configuration files based on a summary list [88, 23] of XML attributes used to bind icons to extract the icon sources associated with the widgets. For UI interactions like clicks, we extract events and their corresponding callback methods, such as onClick. Widgets in Android apps may bind to callback methods either through XML attributes or programmatically through API calls like findViewByID, which links to a unique widget ID. To analyze these bindings comprehensively, we employ GATOR [69, 83, 84], an Android static analysis tool based on Soot [47], to extract widgets and callback methods.
@@
Call Graph Construction. We employ AndroGuard [2], a popular Android static analysis tool, to construct a call graph for each identified callback methods. These graphs elucidate the invocation relationships between methods, with nodes representing methods and edges indicating their interactions. With a callback method as the entry point, each graph not only tracks direct API invocations but also the cascade of method calls they may initiate. As the call graph of a callback method can be considered as a subgraph of the call graph of the entire app [79], our pre-fill pipeline resorts to AndroGuard [2] to construct a complete call graph of the app, and then extract the call graph for each identified callback method. Ultimately, mappings between the widget component and its reachable methods can be obtained.
@@
Android Permission Extraction. This step maps each widget component to its associated potential Android permission requests. Our pre-fill pipeline traverses the call graph of the widget’s callback method to extract all the API calls, and leverages AndroGuard [2] to obtain a mapping from APIs to Android permissions and determine which permissions are requested. This step associates widget components with their corresponding Android permission uses.
@@
Permission-Data Mapping. Our pre-fill pipeline establishes associations between widget components and interaction-driven data practices by mapping Android permissions to relevant data types. Following the approach in prior works [67, 56], we group the dangerous Android permissions into ten different data type categories. For example, a widget associated with Android permission ACCESS_COARSE_LOCATION is considered to be related to data type Location, e.g., the user interaction with this widget leads to the application’s data practice of accessing the user’s location information. Dangerous Android permissions refer to permissions with a protection level of dangerous in the Official Android API Documentation [10].
@@
TPL Detection. Third-party libraries (TPLs) are constitutional in mobile app development. However, insufficient understanding of TPLs used may lead to developers’ non-comprehensive understanding of the privacy behaviors of the application [45, 52, 49, 50, 25]. We adapt the existing state-of-the-art TPL detector, LibScan [78], to extract the TPLs and version information based on code features and sophisticated similarity analysis. It compares each app class with each class in the provided TPL to establish pairwise class correspondence relations, and then computes a confidence score for the presence of the TPL based on the correspondence, eventually reporting the verdict for TPL existence.
@@
4.2 Privacy Notice Analysis
@@
Privacy Policy Segmentation. The objective of this process is to obtain privacy policy segments related to certain data types. For example, segments of data type Location will include all descriptions related to location in the privacy policy, such as whether it will be collected and the purpose of such collection. We adopt the multi-level privacy policies processing strategy in  [62, 82] to perform segment extraction, which takes the app’s privacy policy as input and extract sentences related to each data type through paragraph-level headings classification and sentence-level keyword searching and phrase similarity calculation. We employ prior work [62, 82] for this privacy policy extraction and segmentation process.
Privacy Label Processing. This workflow extracts privacy practice disclosures from the application’s privacy label, e.g. the DSS [5] in Google Play. The privacy label information in the DSS includes the collected data type, the purpose of collection, and whether this collection is optional. Specifically, we record the existence and collection purpose of the data type, as well as whether the granting of this data type is optional for application users. We then map the data types with the data types extracted through the static analysis pipeline discussed in section 4.1. An example of PriBOM is presented in Table 2 of a real mobile application [8] for the UI widget whose ID is “2131296311” and the type is android.view.MenuItem. From the PriBOM, developers can easily obtain that this UI widget is related to permission request, android.permission.ACCESS_COARSE_LOCATION, for coarse positioning and indicates its corresponding data type, Location. In addition, from the Privacy Notice Disclosure section, developers can notice the disclosure in the privacy policy about Coarse Location, “…may collect your geo-location information to optimize user experience…”; and corresponding privacy label Approximate Location information, which is optional for data collection and App functionality, Analytics, Advertising or marketing as its purposes.
5 Human Evaluation
To validate PriBOM’s usefulness, we conducted an online survey to examine the perspectives of people in various software development roles on PriBOM. Ethical approval for this research was secured from our institution’s Institutional Review Board (IRB).
5.1 User Study Design
@@
For the questionnaire, we developed our survey questions based on the specific section design of PriBOM and perspectives from related studies and guidelines [25, 77, 80, 46]. After iterative refinement and pilot studies with 10 participants in total, our final survey contains 26 statements, with which participants rate their agreement levels on a 5-point Likert scale ranging from 1 (strongly disagree) to 5 (strongly agree). Table 4 enumerates the statements. The questions covers six perspectives: (1) General Design of PriBOM, (2) Widget Identifier, (3) Codebase and Permission, (4) TPL, (5) Privacy Notice Disclosure, and (6) Usability and Practicality. Drawing on the pilot study results, we also include four free-text open-ended questions to seek deeper perspectives from participants. The open-ended questions are listed in Table 5.
@@
Table 3: Participants demographics. Each answer is labeled with the count of participant(s) that select it.
@@
Table 4: Our survey results of participants’ agreement on statements related to PriBOM. “Ave.” stands for the average Score, and a higher score indicates stronger agreement. “Distr.” denotes the distribution of the responses (from left to right: strongly disagree to strongly agree).
@@
$\textit{S}_{1}$ [Intuitiveness] The PriBOM’s data fields are logical and promotes ease of understanding. 3.91 [Uncaptioned image]
@@
$\textit{S}_{8}$ [Permission] Listing ‘Permission’ requirements in PriBOM is crucial for transparent disclosure of data access needs and permission requests. 4.09 [Uncaptioned image]
@@
$\textit{S}_{11}$ [Management] PriBOM’s inclusion of third-party library information is crucial for a comprehensive privacy review. 3.85 [Uncaptioned image]
@@
Table 5: The free text open-ended questions in our survey.

### v2

@@ 3. The PriBOM Approach
Figure 3. An overview of PriBOM in the usage scenario. The interaction between app users and UI components triggers data practices. These practices are disclosed to users through privacy notices such as the privacy policy. PriBOM helps development team create accurate privacy notices by documenting privacy information related to specific UI components.
3.1. Motivation of PriBOM
@@
Bill of Materials (BOM). To cater to increasingly complex and diverse software systems, Bill of Materials (BOM) has been introduced as the information inventory with a focus on the transparency of software materials. Proposed in the 2010s (SPD, 2023), SBOM has become a key strategy in response to the emerging risks in software vulnerabilities management, license compliance, and supply chain transparency (Xia et al., 2023b; NTIA, 2019). SBOMs record the material details of the components in developing software, providing transparency, traceability and verifiability to software building and managing process (Stalnaker et al., 2024; Zahan et al., 2023; Mirakhorli et al., 2024). The US Presidential Executive Order (EO) 14028 on Improving the Nation’s Cybersecurity [^3] US Presidential Executive Order 14028 requires companies providing services to the US government to provide SBOMs. Approximately 78% of the organizations worldwide would use SBOMs by 2022 and 88% by 2023 (Stalnaker et al., 2024; Hendrick, 2022). In addition to SBOM, the concept of BOM has been implemented into other forms, including HBOMs (CycloneDX, 2022) and FBOMs (Amy Nelson, 2021; Costin, 2022; Beer, 2022) for hardware and firmware, DataBOMs (Barclay et al., 2019) for datasets, and AIBOMs (Chan, 2017; Xia et al., 2023a) for AI models. Consequently, we adapt a similar design for privacy information and introduce the concept of PriBOM. PriBOM records software privacy information in a structured way and with a focus on UI, pulling different development roles to the same page. Specifically, PriBOM is indexed by the UI widget, including privacy information such as associated privacy permissions, data types, third-party libraries, and corresponding privacy notice disclosures. We will elaborate the design in Section 3.2.
@@
UI Widgets - The Pivot of Mobile Privacy Communication. The UI widgets serve as both visual elements and key components in functionality and data handling, therefore it is the pivot to synergistically connect different roles on privacy communication. First, UI components are fundamental in the software development process, which are not only visual elements but are often directly involved in the app’s functionality and data handling. The transparency regarding what data is collected through specific UI components is one of the information that users want to know (Pan et al., 2024a), especially when malicious design may lead to unwanted data collection and cause harm to users (Brignull, 2010; Chen et al., 2023). Moreover, UI represents the tangible interface where users interact with the application, making it a critical point that is associated with potential privacy issues. When conducting privacy-related tests, interaction with the UI widgets is also a crucial part of executing privacy checks (Wang et al., 2018; Nan et al., 2015). Leveraging UI to link privacy information does inherently focus less on non-UI components, which may lead to gaps in capturing practices that circumvent standard permission protocols (Reardon et al., 2019). Our choice is a deliberate trade-off to prioritize ease of use for all stakeholders over exhaustive documentation, fostering a more accessible approach to privacy management across different roles.
@@
Table 2. Design of PriBOM. One example of filled PriBOM of a real mobile application (Lep, 2024) is presented.
@@ 3.2. Design of PriBOM
  • [Permission] Android utilizes app permission mechanism (Per, 2024) to protect access to both restricted data, e.g. Contacts, and restricted actions, e.g. taking pictures. More specifically, an app must require runtime permissions before it obtains access to additional data or performs actions that may affect the system, other apps or devices. Such requests may be involved in the widget’s behind-the-scene behaviors (Xi et al., 2019). For example, an elliptical coordinate icon in a food delivery app is likely to be accompanied by a request to grant geographic location permissions. Such requests are raised by the Android API calls in the widget callbacks.
@@
  • [Data Type] This field records the specific data type required by corresponding Android Permission, which is also regarded as the data that the widget collects or processes. Following the grouping strategy in previous researches (McConkey and Olukoya, 2023; Rahman et al., 2022) and Official Android API Documentation (Man, 2024), we map the Android Permissions to specific data types. For example, “READ_CONTACTS” and “WRITE_CONTACTS” are categorized into data type “Contacts”. A more detailed discussion is in Section 4.1.
@@
  • [TPL Information] Third-party libraries (TPL) pervasively serve as reusable functional components for mobile apps to improve development efficiency (Wu et al., 2023). For example, the in-app payment service of a shopping app may be supported by TPLs. PriBOM, therefore, records the baseline information of TPLs relevant to the widget, including name, version, latest version, publish date (current version), and publish date (latest version). PriBOM documents the publish date of the currently used and most up-to-date version of TPLs to enable assessing the currency of the version in use and informing about potential updates needed for security or functionality enhancements.
Privacy Notice Disclosure Section. In application development, a significant portion of effort is often allocated to ongoing updates rather than the outset. This principle also applies to privacy notices, given the evolving nature of legal and regulatory requirements. Consequently, PriBOM effectively manages two common forms of privacy notices, e.g. privacy policies and privacy labels, by linking widget components directly to their respective disclosures, ensuring continuous alignment between app data practices and legal requirements throughout the development lifecycle.
@@
  • [Privacy Label Declaration] Before publishing apps on markets such as Google Play, developers must declare their app’s privacy practices, e.g. data collection and handling, in a privacy nutrition label form. However, developers often struggle to create authoritative and accurate privacy labels (Li et al., 2022; Balebako and Cranor, 2014; Balebako et al., 2014). Including privacy labels’ data practice declarations that match the data type in PriBOM can support developers and legal teams in aligning privacy disclosure with actual app behaviors.
Figure 4. The pipeline of our static analysis module.
3.3. Benifits of PriBOM
@@
[Benefit-1] Transparency. Lacking team-level and organizational support is considered a challenge faced by developers (Li et al., 2018, 2022). Therefore, to address the privacy ambiguity caused by the lack of team support and assistant tools, the necessity of a recording and communication tool focused on privacy information has become more prominent. PriBOM can serve as a central communication tool, bridging the informational gap between developers of different roles from various teams. By providing a unified view of privacy practices at the widget level, PriBOM can help all developers have a consistent understanding of potential privacy implications, fostering informed decision-making and cohesive development strategies. Through PriBOM, transparency can be achieved as every practitioner, regardless of their role, gains clear insights into privacy aspects.
[Benefit-2] Privacy Traceability and Trackability. PriBOM connects the frontend UI widget, the backend privacy practices in source code, and the app’s privacy notices, helping developers gain traceability and trackability on this privacy chain. Similar to terminology in software supply chain (Kelepouris et al., 2007), both terms in the context of PriBOM with ‘where-from’ and ‘where-used’ (Petroff and Hill, 1991) directions respectively are as follows:
  • Tracing. Given any point in this privacy chain, PriBOM enables developers to work backward to query privacy-relevant information. For example, if a user reports privacy issues in interacting with certain UI widgets, then the development team can easily track back and locate the problematic code files (Currently, a privacy breach takes an average of 287 days to completely resolve (sim, 2024)).
@@ 4. Pre-fill of PriBOM
In this section, we present a pre-fill implementation of PriBOM, and describe details regarding the pipeline. Our pre-fill implementation is built upon existing tools (Rountev and Yan, 2014; Yang et al., 2015, 2018; And, 2024; Wu et al., 2023; Pan et al., 2024a). Notably, although our design tailors to Android apps, the concept of PriBOM can be easily adapted on iOS apps with corresponding modifications.
4.1. Static Analysis
@@
Following Android analysis approaches in previous icon-behavior researches (Xi et al., 2019; Zhao et al., 2021; Malviya et al., 2023), our static analysis-based pre-fill implementation consists of four steps: (1) Widget and Callback Method Extraction; (2) Call Graph Construction; (3) Android Permission Extraction; and (4) Permission-Data Mapping. Figure 4 provides an overview of the pipeline.
@@
Widget and Callback Method Extraction. In the first step, the implementation pipeline takes the Android application package (APK) file as input, and leverages JADX (jad, 2024) to disassemble it. The purpose of this step is to extract widget components, associated sources such as icons, and corresponding callback methods that respond to UI interactions. Icons (such as PNG or JPG files) are commonly bound with UI widgets in the layout configuration file. For example, XML attribute android: background is used to set a background image for a widget component. Therefore, we parse the layout configuration files based on a summary list (Zhao et al., 2021; Avdiienko et al., 2017) of XML attributes used to bind icons to extract the icon sources associated with the widgets. For UI interactions like clicks, we extract events and their corresponding callback methods, such as onClick. Widgets in Android apps may bind to callback methods either through XML attributes or programmatically through API calls like findViewByID, which links to a unique widget ID. To analyze these bindings comprehensively, we employ GATOR (Rountev and Yan, 2014; Yang et al., 2015, 2018), an Android static analysis tool based on Soot (Lam et al., 2011), to extract widgets and callback methods.
@@
Call Graph Construction. We employ AndroGuard (And, 2024), a popular Android static analysis tool, to construct a call graph for each identified callback methods. These graphs elucidate the invocation relationships between methods, with nodes representing methods and edges indicating their interactions. With a callback method as the entry point, each graph not only tracks direct API invocations but also the cascade of method calls they may initiate. As the call graph of a callback method can be considered as a subgraph of the call graph of the entire app (Xi et al., 2019), our pre-fill pipeline resorts to AndroGuard (And, 2024) to construct a complete call graph of the app, and then extract the call graph for each identified callback method. Ultimately, mappings between the widget component and its reachable methods can be obtained.
@@
Android Permission Extraction. This step maps each widget component to its associated potential Android permission requests. Our pre-fill pipeline traverses the call graph of the widget’s callback method to extract all the API calls, and leverages AndroGuard (And, 2024) to obtain a mapping from APIs to Android permissions and determine which permissions are requested. This step associates widget components with their corresponding Android permission uses.
@@
Permission-Data Mapping. Our pre-fill pipeline establishes associations between widget components and interaction-driven data practices by mapping Android permissions to relevant data types. Following the approach in prior works (Rahman et al., 2022; McConkey and Olukoya, 2023), we group the dangerous Android permissions into ten different data type categories. For example, a widget associated with Android permission ACCESS_COARSE_LOCATION is considered to be related to data type Location, e.g., the user interaction with this widget leads to the application’s data practice of accessing the user’s location information. Dangerous Android permissions refer to permissions with a protection level of dangerous in the Official Android API Documentation (Man, 2024).
@@
TPL Detection. Third-party libraries (TPLs) are constitutional in mobile app development. However, insufficient understanding of TPLs used may lead to developers’ non-comprehensive understanding of the privacy behaviors of the application (Khandelwal et al., 2023; Li et al., 2022, 2018, 2021a; Balebako et al., 2014). We adapt the existing state-of-the-art TPL detector, LibScan (Wu et al., 2023), to extract the TPLs and version information based on code features and sophisticated similarity analysis. It establishes pairwise class correspondences between app and TPL classes, computes confidence scores, and determines TPL presence.
@@
4.2. Privacy Notice Analysis
@@
Privacy Policy Segmentation. The objective of this process is to obtain privacy policy segments related to certain data types. For example, segments of data type Location will include all descriptions related to location in the privacy policy, such as whether it will be collected and the purpose of such collection. We adopt the multi-level privacy policies processing strategy in  (Pan et al., 2024a; Xie et al., 2022) to perform segment extraction, which takes the app’s privacy policy as input and extract sentences related to each data type through paragraph-level headings classification and sentence-level keyword searching and phrase similarity calculation. We employ prior work (Pan et al., 2024a; Xie et al., 2022) for this privacy policy extraction and segmentation process.
Privacy Label Processing. This workflow extracts privacy practice disclosures from the application’s privacy label, e.g. the DSS (Dat, 2024) in Google Play. The privacy label information in the DSS includes the collected data type, the purpose of collection, and whether this collection is optional. Specifically, we record the existence and collection purpose of the data type, as well as whether the granting of this data type is optional for application users. We then map the data types with the data types extracted through the static analysis pipeline discussed in section 4.1. An example of PriBOM is presented in Table 2 of a real mobile application (Lep, 2024) for the UI widget whose ID is “2131296311” and the type is android.view.MenuItem. From the PriBOM, developers can easily obtain that this UI widget is related to permission request, android.permission.ACCESS_COARSE_LOCATION, for coarse positioning and indicates its corresponding data type, Location. In addition, from the Privacy Notice Disclosure section, developers can notice the disclosure in the privacy policy about Coarse Location, “…may collect your geo-location information to optimize user experience…”; and corresponding privacy label Approximate Location information, which is optional for data collection and App functionality, Analytics, Advertising or marketing as its purposes.
5. Human Evaluation
To explore PriBOM’s usefulness and prompt further requirements, we conducted a totally anonymous online survey to examine the perspectives of people in various software development roles on PriBOM. Ethical approval for this research was secured from our institution’s Institutional Review Board (IRB). For more information, please refer to Section 5.3.
5.1. User Study Design
@@
For the questionnaire, we developed our survey questions based on the specific section design of PriBOM and perspectives from related studies and guidelines (Balebako et al., 2014; Weir et al., 2020; Xia et al., 2023a; Kitchenham and Pfleeger, 2008). After iterative refinement and pilot studies with 10 participants in total, our final survey contains 26 statements, with which participants rate their agreement levels on a 5-point Likert scale ranging from 1 (strongly disagree) to 5 (strongly agree). Table 4 enumerates the statements. The questions covers six perspectives: (1) General Design of PriBOM, (2) Widget Identifier, (3) Codebase and Permission, (4) TPL, (5) Privacy Notice Disclosure, and (6) Usability and Practicality. Drawing on the pilot study results, we also include four free-text open-ended questions to seek deeper perspectives from participants. The open-ended questions are listed in Table 5.
@@
Table 3. Participants demographics. Each answer is labeled with the count of participant(s) that select it. “N.Am.” stands for “North America”.
@@
Table 4. Our survey results of participants’ agreement on statements related to PriBOM. “Ave.” stands for the average Score, and a higher score indicates stronger agreement. “Distr.” denotes the distribution of the responses (from left to right: strongly disagree to strongly agree).
@@
$\textit{S}_{1}$ [Intuitiveness] The data fields in PriBOM appear logically structured and are understandable. 3.91 [Uncaptioned image]
@@
$\textit{S}_{8}$ [Permission] Listing ‘Permission’ requirements in PriBOM may increase transparency regarding data access needs and permission requests. 4.09 [Uncaptioned image]
@@
$\textit{S}_{11}$ [Management] Including third-party library information in PriBOM may support a more thorough privacy review. 3.85 [Uncaptioned image]
@@
5.2. Participant Recruitment
Following the participant recruitment processes in prior works in Usable Privacy and Security field (Xia et al., 2023a; Lin et al., 2023; Pan et al., 2024a), we published the survey via Qualtrics (Qua, 2024), and recruited participants through Prolific (Pro, 2024), a commonly used recruitment platform of sustainably gathering survey responses at scale. We use the participant screener in Prolific to recruit participants that meet our requirement discussed in Section 5.1.
We consider responses that are uncompleted or with a completion time of less than 2 minutes to be invalid. In total, we received 150 valid survey responses from participants with diverse backgrounds. The median time to complete the survey is 14 minutes and 33 seconds. Our participants span across 35 countries include 94 males and 55 females, with one participant preferring not to say, which conforms the reality that there are more male practitioners in software industry. We also include 3 demographic questions in the survey questionnaire to understand the roles of participants in software development, years of experience, and the size of the team they work in. An overview of the demographic information is presented in Table 3. The role with the highest proportion is junior developer (39%), while other participants reported to be UI designers (15%), project managers (14%), senior developers (13%), legal team members (9%), and other roles such as CEO. The team size among participants is also diverse. The most common team size is less than 10 people (59%), while there are also 9% of participants working in a team with more than 50 people. We believe our participants’ backgrounds are sufficiently diverse for our survey. For ethics consideration, please refer to Section 5.3.
5.3. Ethics Considerations
We recruited participants through Prolific (Pro, 2024), a highly regarded online platform for researchers to recruit participants for UPS studies. We only collect basic and non-identifiable demographic information. Participants receive monetary rewards at a rate of £7.49 per hour, as recommended by Prolific. We ensure the risk of participating in this research is minimum, and no disturbing content is distributed in the survey. A consent and withdrawal information sheet was presented before the survey to inform participants of their rights, including the ability to withdraw at any point before submitting by exiting the survey page. Participants give their consent by proceeding with the survey. All information provided by participants is treated confidentially and we will not share the collected data to any third parties.
Table 5. The free text open-ended questions in our survey.